OpenStudy (anonymous):

you are debating whether to buy a new computer for 1,360.00 with a discount of 10% and sales tax of 6.5% or a refurbrished computer with the same equipment for 1,095.00 and sales tax of 5.75% if a savings account earns 3.75% APR how much do you really save with a refurbrised computer if you put the difference into savings accountsfor a year and calculate using the simple interest formula

OpenStudy (ranga):

10% (or 0.1) discount on the new computer. So the discounted price will be 0.9 times the price or 1360 * 0.9 = $1,224 Sales tax is 6.5%. Total cost = $1224 + 1224 * 0.065 = $1303.56 Refurbished computer cost = $1,095 * 1.0575 = $1157.96 Savings = $1303.56 - $1157.96 = $145.60 Invest the savings of $145.60 at 3.75% APR. Interest earned = P * r * t = 145.60 * 0.0375 * 1 = $5.46 Total savings = $145.60 + $5.46 = $151.06

OpenStudy (anonymous):

thx :)

OpenStudy (ranga):

yw.